# Exploiting Completeness Perception with Diffusion Transformer for Unified 3D MRI Synthesis

### v1

@@ 4.3 Ablation Study
Effect of Prompt Tokens. Table 4 presents the synthesis performance associated with each learned prompt token and mask code. The mask codes are generated as one-hot vectors whose dimension corresponds to the number of modalities or slices, indicating the corresponding missing positions. Notably, we employ an identical injection strategy for all prompts and masks to ensure a fair comparison. As shown, all individual prompt tokens and their combinations improve model performance, surpassing guidance based on binary mask codes. These findings strongly highlight the effectiveness and rationale of our completeness perception scheme. The incompleteness positioning task yields the most effective prompts, likely due to its explicit emphasis on positioning missing sections along the modality/slice dimension, thus increasing the sensitivity to subtle structural variations. To further explore the potential of our learned prompts, we incorporate them into prior baselines as replacements for mask codes. As illustrated in Table 13, this substitution consistently improves their performance, demonstrating the utility and generalizability of our completeness-aware prompts, which can serve as a flexible paradigm for improving emerging MRI synthesis methods.
@@
Choice of Prompt Injection Position. This experiment aims to justify our choice of injecting prompts exclusively into the modal and spatial blocks for brain and cardiac tasks, respectively. As illustrated in Table 14, injecting prompts into the modal/spatial block yields the best overall performance, whereas injecting into spatial/planar blocks or both blocks results in inferior outcomes. This can be attributed to the fact that aligning the prompt with the block that explicitly models the task’s primary dependency, namely modality fusion for brain and through-plane continuity for cardiac, maximizes the efficacy of conditional signals.
Impact of 3D MRI Diffusion Transformer. The quantitative evaluation of MDiT3D and existing diffusion networks [39, 34, 30] is presented in Table 7. The results illustrate that MDiT3D consistently achieves superior performance, confirming that the specialized architectural designs effectively enhance the quality of 3D MRI synthesis.
@@ 4.4 Visualization Analysis
Figure 6: Visualization of salient regions on BraTS dataset.
Figure 7: t-SNE visualization on the BraTS dataset of (a) The count-focused prompt $\mathbf{p}^{d}$ colored by missing-number classes; and (b) The identity-focused prompt $\mathbf{p}^{p}$ colored by missing modality.
Figure 8: Visualization of attention maps from individual modal blocks on BraTS dataset under a missing configuration of [1, 0, 0, 0], where the first modality is absent.
@@
Salient Regions. To better understand the learning procedure of pretext tasks, we visualize their activation maps using GradCAM [42]. As shown in Fig. 6, the four examples reveal strong correlations between salient regions and modality-discriminative features. Specifically, these patterns mirror each task’s objective: Task 1 and Task 3 must assess global consistency, so they rely on coarse anatomical layout (gray and white matter) that summarize the volume. In contrast, Task 2 needs to pinpoint where/which is missing, so it keys on high-frequency, modality-specific cues (tumors, lesions, and white matter hyperintensities in FLAIR) to localize. The results emphasize that our pretext tasks effectively capture modality-specific properties, understanding the MRI context and relationships.
@@
Prompt Distribution. We implement the t-SNE visualization of learned prompts on BraTS and color them by ground-truth incompleteness labels. As depicted in Fig. 7, the count-focused prompt $\mathbf{p}^{d}$ forms compact, well-separated clusters aligned with the missing number classes corresponding to each missing state (1/2/3 absent modalities), while the identity-focused prompt $\mathbf{p}^{p}$ produces modality-specific clusters (T1, T1ce, T2, FLAIR) with clear boundaries, highlighting that the two prompts capture complementary aspects of incompleteness.
@@
Attention Maps. We investigate the contribution of prompts by visualizing the attention maps of each modal block. As shown in Fig. 8, our prompts guide the attention mechanism to progressively focus on the actual missing elements (the first row and column) with the increase of block layers. In comparison, mask codes lack sufficient informativeness and fail to provide nuanced guidance that aligns with the true missing state.
@@ 5 Conclusion
This work presents CoPeDiT, a unified model for 3D brain and cardiac MRI synthesis that explores completeness perception. We demonstrate that enabling the model to autonomously infer missing state, rather than relying on externally pre-defined masks, can provide more discriminative and informative guidance. To this end, we equip our tokenizer with completeness perception capability through elaborated pretext tasks. MDiT3D is then developed to utilize the learned prompt tokens as guidance for 3D MRI generation. Experimental evaluation validates the superiority of CoPeDiT in terms of accuracy, generalizability, robustness, and adaptability across diverse missing scenarios. Furthermore, CoPeDiT enables mask-free synthesis through perceiving completeness state, enhancing flexibility and applicability in practical clinical settings.
Supplementary Material
@@ Appendix A Datasets
BraTS [3] T1, T1ce, T2, FLAIR 1251 1000 251
IXI [5] T1, T2, PD 577 462 115
@@
UKBB [35] - 31350 25080 6270
MESA [57] - 298 238 60
ACDC [4] - 300 240 60
MSCMR [59] - 300 240 60
@@ B.1 Data Preprocessing
Brain MRI data. Following [24, 29], we use 90 and 80 middle axial slices for BraTS and IXI datasets, respectively. These slices are further cropped to a size of $192\times 192$ from the central region. Ultimately, all volumes are resized to a fixed size of $192\times 192\times 64$ to serve as model input.
@@ B.2 Architecture of Prompt Encoders and Projection Heads
Notably, in our framework, binary mask codes are only used in the pretraining of CoPeVAE, where we synthetically remove modalities/slices and supervise the pretext tasks with known missing patterns. Once CoPeVAE is trained, we freeze its parameters and use it as a completeness-aware tokenizer: given any incomplete MRI with an arbitrary missing pattern, CoPeVAE directly infers the corresponding completeness prompts $\mathbf{p}=\mathbf{p}^{d}\,\|\,\mathbf{p}^{p}\,\|\,\mathbf{p}^{s}$ from the observed data, without requiring explicit mask codes as input. During both diffusion model training and inference, the diffusion backbone receives only the latent representations and these learned prompts. The original binary masks that were used to generate synthetic missingness are not provided to the diffusion model. In this sense, our synthesis process is mask-free: CoPeDiT no longer depends on hand-crafted or externally supplied mask codes at the generation stage, but instead relies entirely on the learned completeness prompts produced by the frozen CoPeVAE.
@@ B.3 Hyperparameter Setups
CoPeVAE. The detailed hyperparameter setup of CoPeVAE is provided in Table 10. Built upon VQVAE [47] and VQGAN [10], our model employs a codebook containing 8192 codes with the latent dimensionality of 8. The values of $\tau$ and $\lambda$ are empirically set, as they have only a slight impact on model performance. The Adam optimizer is applied with a warmup cosine learning rate schedule. The training steps of CoPeVAE-B and CoPeVAE-C are 400k and 100k, respectively.
Figure 9: Qualitative results of comparison on IXI dataset. The results are depicted in the axial (top), sagittal (middle) and coronal (bottom) views of the 3D MRI volume.
@@
Training objective x-prediction x-prediction
@@
MDiT3D. For MDiT3D, we carefully design the hyperparameters to balance dataset size and model capacity, as summarized in Table 11. Following DiT [34], we report the experimental results using the exponential moving average (EMA) with a decay rate of 0.9999. During training, we set the timestep to 500 with linearly scaled noise levels ranging from 0.0015 to 0.0195. MDiT3D is trained for 100k iterations using the AdamW optimizer and a warmup cosine learning rate schedule. During inference, the DDIM sampler is applied with sample steps of 200 and 250 for MDiT3D-B and MDiT3D-C, respectively.
@@ C.1 Qualitative Results
Figure 10: Sensitivity of CoPeDiT to prompt corruption on BraTS dataset.
Table 12: Experimental results of tumor segmentation experiments on BraTS dataset.
[…]
WT TC ET AVG
Missing 86.08 84.67 81.59 84.11
GAN-based Methods
MMGAN [43] 89.35 88.14 87.73 88.41
MMT [24] 90.43 88.37 86.92 88.57
Hyper-GAE [53] 88.72 86.54 85.37 86.88
Diffusion Model-based Methods
LDM [39] 87.86 85.91 84.19 85.99
ControlNet [58] 88.27 87.05 85.23 86.85
M2DN [29] 91.28 90.09 88.20 89.86
CoPeDiT 91.35 90.41 88.94 90.23
C.2 Tumor Segmentation
In this section, we evaluate the clinical utility of our synthesized MRIs through downstream tumor segmentation experiments on the BraTS dataset. Specifically, following [24, 29], we train a multi-modal U-Net [40] on the four MRI modalities, where one modality is withheld and replaced by the synthesized image generated by our CoPeDiT or competing baselines. Additionally, we implement a U-Net that takes only the three available modalities as input for comparison. We report the average Dice scores for whole tumor (WT), tumor core (TC), and enhancing tumor (ET), as summarized in Table 12. As shown, all methods improve the segmentation accuracy over the ’Missing’ baseline that only uses the three available modalities, confirming the benefit of recovering the missing contrast. Among them, CoPeDiT achieves the best performance on all three tumor subregions, leading to the highest average Dice of 90.23%. These results demonstrate that our synthesized MRIs are not only visually plausible but also provide more informative inputs for downstream clinical tasks such as tumor segmentation.
@@ C.2 Tumor Segmentation
w/o $\mathbf{p}^{d}$ 23.56 0.720 35.93 23.04 0.708 48.71
w/o $\mathbf{p}^{p}$ 22.81 0.697 50.70 21.99 0.692 63.16
w/o $\mathbf{p}^{s}$ 24.17 0.724 31.89 23.56 0.710 39.87
w/o Prompts 22.48 0.696 57.64 21.67 0.683 81.49
w/ Mask Codes 23.35 0.711 44.03 22.84 0.702 56.72
CoPeDiT 24.34 0.732 25.84 23.92 0.721 32.53
@@
MMT [24] 22.64 0.698 53.60
+ Prompts (ours) 23.19 (+0.55) 0.707 (+0.009) 46.13 (–7.47)
Hyper-GAE [53] 22.12 0.682 72.62
+ Prompts (ours) 22.46 (+0.34) 0.694 (+0.012) 59.43 (–13.19)
M2DN [29] 23.47 0.715 42.52
+ Prompts (ours) 23.84 (+0.37) 0.726 (+0.011) 33.79 (–8.73)
C.3 Prompt Accuracy Sensitivity
To assess how sensitive CoPeDiT is to the correctness of completeness prompts, we conduct a controlled prompt corruption study on the BraTS dataset under the number of missing modalities of 1. Starting from a trained CoPeDiT and CoPeVAE, we consider the degree prompt $\mathbf{p}^{d}$ (number/length of missing content), the position prompt $\mathbf{p}^{p}$ (location of missing regions), and the semantic prompt $\mathbf{p}^{s}$. For each prompt type, we define a corruption rate $r\in\{0,0.25,0.5,0.75,1.0\}$. Given a rate $r$, we randomly select an $r$-fraction of validation samples and replace the corresponding prompt with an incorrect one (sampled from other cases), while keeping all network weights and the other prompts unchanged. We also include an ’All’ setting where all three prompts are corrupted simultaneously. As illustrated in Fig. 10, increasing the corruption rate consistently degrades synthesis quality in terms of all three metrics. This demonstrates that CoPeDiT relies on the accuracy of completeness prompts. Among the three types, corrupting the position prompt $\mathbf{p}^{p}$ causes the largest synthesis performance drop, emphasizing that precise localization of missing regions is particularly critical. Corrupting the degree prompt $\mathbf{p}^{d}$ also noticeably harms performance, whereas the semantic prompt $\mathbf{p}^{s}$ is somewhat more robust but still shows a clear degradation trend. When all prompts are corrupted, performance deteriorates the most across all metrics, confirming that the three prompts provide complementary and influential guidance for high-fidelity synthesis.
C.4 Ablation Study

### v2

@@ 4.3 Ablation Study
Effect of Prompt Tokens. Table 4 compares learned prompt tokens against one-hot mask codes. Under identical injection strategies, all individual prompts and their combinations consistently outperform binary masks, strongly validating our completeness perception approach. Specifically, the positioning prompt ($p^{p}$) proves most effective; its removal triggers the sharpest performance drop, likely because explicitly locating missing sections heightens sensitivity to subtle structural variations. To further explore the potential of our learned prompts, we incorporate them into prior baselines as replacements for mask codes. As illustrated in Table 6, this substitution consistently boosts their performance, such as yielding a 0.78 dB PSNR gain for M2DN, demonstrating the strong plug-and-play utility and generalizability of our method.
@@
Choice of Prompt Injection Position. We aim to justify our choice of injecting prompts exclusively into the modal and spatial blocks for brain and cardiac tasks, respectively. As shown in Table 7, injecting prompts into the modal/spatial block yields the best performance, whereas injecting into spatial/planar blocks or both blocks results in inferior outcomes. This can be attributed to the fact that aligning the prompt with the block that explicitly models the task’s primary dependency, namely modality fusion for brain and through-plane continuity for cardiac, maximizes the efficacy of conditional signals.
Impact of 3D MRI Diffusion Transformer. The evaluation of MDiT3D and existing diffusers Rombach et al. (2022); Peebles & Xie (2023); Mo et al. (2023) is presented in Table 8. MDiT3D achieves superior performance, notably outperforming vanilla 3D DiTs by over 1.4 dB in PSNR on the BraTS dataset. This confirms that our task-driven design, which couples alternating blocks with targeted prompt injection to explicitly model inter-modal relationships and through-plane continuity, effectively improves synthesis.
Table 9: Results of tumor segmentation experiments on the BraTS dataset.
[…]
WT TC ET AVG
Missing 86.08 84.67 81.59 84.11
GAN-based Methods
MMGAN Sharma & Hamarneh (2020) 89.35 88.14 87.73 88.41
MMT Liu et al. (2023) 90.43 88.37 86.92 88.57
Hyper-GAE Yang et al. (2023) 88.72 86.54 85.37 86.88
Diffusion Model-based Methods
LDM Rombach et al. (2022) 87.86 85.91 84.19 85.99
ControlNet Zhang et al. (2023) 88.27 87.05 85.23 86.85
M2DN Meng et al. (2024) 91.28 90.09 88.20 89.86
DiffM4RI Ye et al. (2026) 90.04 89.23 87.68 88.72
CoPeDiT 91.35 90.41 88.94 90.23
Figure 6: Visualization of salient regions on the BraTS dataset.
4.4 Tumor Segmentation
To evaluate clinical utility, we conduct downstream tumor segmentation on the BraTS dataset. Following Liu et al. (2023); Meng et al. (2024), we train a multi-modal U-Net Ronneberger et al. (2015) using three available modalities alongside one modality synthesized by CoPeDiT or baselines. We also evaluate a "Missing" baseline trained solely on the three available modalities. As shown in Table 9, all synthesis methods improve upon the "Missing" baseline in terms of average Dice scores for whole tumor (WT), tumor core (TC), and enhancing tumor (ET). Notably, CoPeDiT consistently outperforms competing methods across all subregions, achieving the highest average Dice of 90.23%. These results confirm that our synthesized MRIs provide highly informative and clinically valuable inputs for downstream tasks.
4.5 Computational Efficiency Analysis
Table 10: The computational cost and wall-clock time comparison of CoPeVAE on the BraTS dataset.
Param. (M) Flops (G) Epoch Time (s) Total Time (GPU hours) BraTS (Avg)
[…]
w/o Task 1 136.77 21167.61 331.3 561.2 27.03 455.73
w/o Task 2 136.55 21171.83 332.9 563.9 26.47 468.13
w/o Task 3 136.74 21171.83 330.7 560.1 27.23 401.79
CoPeVAE 137.63 21184.57 336.7 570.3 28.10 288.49
Table 11: The inference cost comparison of CoPeDiT on the BraTS dataset.
Inference Latency (s) VRAM (G)
LDM Rombach et al. (2022) 1.38 0.429
M2DN Meng et al. (2024) 4.67 0.819
DiffM4RI Ye et al. (2026) 2.16 0.584
DiT-3D Mo et al. (2023) 1.65 0.614
CoPeDiT 1.73 0.626
Training Compute and Wall-Clock Cost. As shown in Table 4.5, our CoPeVAE adds minimal overhead in parameters, FLOPs, and training time compared to variants. This modest cost yields superior synthesis quality and spatial consistency, proving a highly efficient complexity-gain trade-off for our pretext tasks.
Inference Cost. To assess practical applicability, Table 4.5 reports end-to-end latency and VRAM usage per volume under standardized settings (batch size = 1, single A100 GPU, DDIM Song et al. (2021) 200 steps, mixed precision). CoPeDiT remains efficient, with resource demands comparable to standard models and markedly lower than heavier baselines, supporting its suitability for clinical deployment.
Figure 7: Qualitative analysis of the learned prompts on the BraTS dataset. (a) t-SNE projection of the count-focused prompt $\mathbf{p}^{d}$, showing clear clustering by the number of missing modalities. (b) t-SNE projection of the identity-focused prompt $\mathbf{p}^{p}$, revealing distinct separation by missing modality types. (c) Visualization of modal block attention maps under a missing configuration of [1, 0, 0, 0].
@@ 4.6 Visualization Analysis
Salient Regions. To understand the learning procedure of pretext tasks, we visualize their activation maps using GradCAM Selvaraju et al. (2017). As shown in Fig. 6, the examples reveal strong correlations between salient regions and modality-discriminative features. These patterns mirror each task’s objective: Task 1 and Task 3 must assess global consistency, so they rely on coarse anatomical layout (gray and white matter) that summarize the volume. In contrast, Task 2 needs to pinpoint where/which is missing, so it keys on high-frequency, modality-specific cues (tumors, lesions, and white matter hyperintensities in FLAIR). The results emphasize that our pretext tasks capture modality-specific properties and improve the model’s understanding of MRI context and inter-modality relationships.
@@
Prompt Distribution. We implement the t-SNE visualization of learned prompts on BraTS and color them by ground-truth incompleteness labels. As depicted in Figs. 7 and 7, the count-focused prompt $\mathbf{p}^{d}$ forms compact, well-separated clusters aligned with the missing number classes corresponding to each missing state (1/2/3 absent modalities). Meanwhile, the identity-focused prompt $\mathbf{p}^{p}$ produces modality-specific clusters (T1, T1ce, T2, FLAIR) with distinct boundaries. This clear separation visually confirms that our pretext tasks effectively drive the tokenizer to learn highly discriminative features. Together, they demonstrate that our learned prompt tokens capture explicitly decoupled and complementary semantic priors. By avoiding feature entanglement, these prompts successfully provide the network with nuanced, multi-granular guidance for reliable synthesis.
@@
Attention Maps. We visualize the attention maps of each modal block under a specific missing configuration (e.g., [1, 0, 0, 0]). As shown in Fig. 7, our learned prompts actively guide the attention mechanism to progressively focus on the actual missing elements (the first row and column) as the block depth increases. This concentrated focus facilitates precise, layer-wise feature aggregation for the absent modality. In comparison, explicit mask codes lack sufficient informativeness, yielding diffuse attention weights that fail to provide the nuanced, dynamic guidance necessary to align with the true missing state.
Figure 8: Quantitative analysis of completeness-aware prompts on the BraTS dataset. (a) Sensitivity of CoPeDiT to increasing prompt corruption rates. Synthesis quality degrades monotonically, particularly when perturbing the position prompt $\mathbf{p}^{p}$. (b) Prompt intervention matrix illustrating the PSNR drop under structured perturbations. Misleading signals (e.g., random, type-swap) and the simultaneous perturbation of all prompts cause the most severe performance degradation.
4.7 Analyzing Completeness-Aware Prompts
To quantitatively validate the effectiveness and rationale of our learned prompts, we conduct comprehensive sensitivity and intervention analyses.
Prompt Accuracy Sensitivity. To assess CoPeDiT’s reliance on prompt accuracy, we randomly replace the degree ($\mathbf{p}^{d}$), position ($\mathbf{p}^{p}$), semantic ($\mathbf{p}^{s}$), or "All" prompts with incorrect ones for an $r$-fraction ($r\in[0,1.0]$) of BraTS validation samples. Fig. 8 demonstrates that synthesis quality degrades consistently as the corruption rate $r$ increases. Corrupting all prompts simultaneously yields the worst outcomes, confirming their complementary and essential roles.
Prompt Intervention. We further apply structured interventions by replacing prompt components with zeroed, random, shuffled, or type-swapped variants. As Fig. 8 illustrates, misleading signals (random/type-swapped) degrade performance more severely than missing or mildly perturbed guidance (zeroed/shuffled). Most notably, across both sensitivity and intervention experiments, perturbing $\mathbf{p}^{p}$ consistently triggers the sharpest performance drops. This combined evidence robustly demonstrates that precise missing region localization ($\mathbf{p}^{p}$) is the most crucial prompt for guiding high-fidelity 3D MRI synthesis.
@@ 5 Conclusion
This work presents CoPeDiT, a unified model for 3D brain and cardiac MRI synthesis that explores completeness perception. We demonstrate that enabling the model to autonomously infer the missing state, rather than relying on externally pre-defined masks, can provide more discriminative and informative guidance. To this end, we equip our tokenizer with completeness perception capability through carefully designed pretext tasks. MDiT3D is then developed to utilize the learned prompt tokens as guidance for 3D MRI generation. Extensive evaluations validate CoPeDiT’s remarkable accuracy and robustness across diverse scenarios, highlighting its potential for practical clinical deployment.
Limitations. While highly effective, CoPeDiT requires a fixed number of modalities during training and may lose some fine high-frequency details due to latent space compression. Future work will focus on modality-agnostic tokenizers and exploring pixel-space diffusion refinement.
Appendix
@@ Appendix A Datasets
BraTS Baid et al. (2021) T1, T1ce, T2, FLAIR 1251 1000 251
IXI Brain Development Project (2025) T1, T2, PD 577 462 115
@@
UKBB Petersen et al. (2016) - 31350 25080 6270
MESA Zhang et al. (2018) - 298 238 60
ACDC Bernard et al. (2018) - 300 240 60
MSCMR Zhuang et al. (2022) - 300 240 60
@@ B.1 Data Preprocessing
Brain MRI data. Following Liu et al. (2023); Meng et al. (2024), we use 90 and 80 middle axial slices for BraTS and IXI datasets, respectively. These slices are further cropped to a size of $192\times 192$ from the central region. Ultimately, all volumes are resized to a fixed size of $192\times 192\times 64$ to serve as model input.
@@ B.2 Architecture of Prompt Encoders and Projection Heads
Notably, in our framework, binary mask codes are only used in the pretraining of CoPeVAE, where we synthetically remove modalities/slices and supervise the pretext tasks with known missing patterns. Once CoPeVAE is trained, we freeze its parameters and use it as a completeness-aware tokenizer: given any incomplete MRI with an arbitrary missing pattern, CoPeVAE directly infers the corresponding completeness prompts $\mathbf{p}=\mathbf{p}^{d}\,\|\,\mathbf{p}^{p}\,\|\,\mathbf{p}^{s}$ from the observed data, without requiring explicit mask codes as input. During both diffusion model training and inference, the diffusion backbone receives only the latent representations and these learned prompts. The original binary masks that were used to generate synthetic missingness are not provided to the diffusion model. In this sense, CoPeDiT no longer depends on hand-crafted or externally supplied mask codes at the generation stage, but instead relies entirely on the learned completeness prompts produced by the frozen CoPeVAE.
@@ B.3 Hyperparameter Setups
CoPeVAE. The detailed hyperparameter setup of CoPeVAE is provided in Table 14. Built upon VQVAE van den Oord et al. (2017) and VQGAN Esser et al. (2021), our model employs a codebook containing 8192 codes with the latent dimensionality of 8. The values of $\tau$ and $\lambda$ are empirically set, as they have only a slight impact on model performance. The Adam optimizer is applied with a warmup cosine learning rate schedule. The training steps of CoPeVAE-B and CoPeVAE-C are 400k and 100k, respectively.
@@
Training objective $\mathbf{x}_{0}$-prediction $\mathbf{x}_{0}$-prediction
@@
MDiT3D. For MDiT3D, we carefully design the hyperparameters to balance dataset size and model capacity, as summarized in Table 15. Following DiT Peebles & Xie (2023), we report the experimental results using the exponential moving average (EMA) with a decay rate of 0.9999. During training, we set the timestep to 500 with linearly scaled noise levels ranging from 0.0015 to 0.0195. MDiT3D is trained for 100k iterations using the AdamW optimizer and a warmup cosine learning rate schedule. During inference, the DDIM sampler Song et al. (2021) is applied with sample steps of 200 and 250 for MDiT3D-B and MDiT3D-C, respectively.
@@ C.1 Qualitative Results
w/o $\mathbf{p}^{d}$ 23.56 0.720 35.93 724.18 23.04 0.708 48.71 917.76
w/o $\mathbf{p}^{p}$ 22.81 0.697 50.70 1267.70 21.99 0.692 63.16 1281.31
w/o $\mathbf{p}^{s}$ 24.17 0.724 31.89 685.13 23.56 0.710 39.87 876.94
w/o Prompts 22.48 0.696 57.64 1315.59 21.67 0.683 81.49 1542.24
w/ Mask Codes 23.35 0.711 44.03 783.46 22.84 0.702 56.72 1067.16
CoPeDiT 24.34 0.732 25.84 569.22 23.92 0.721 32.53 718.54
@@
MMT Liu et al. (2023) 22.64 0.698 53.60 1329.25
+ Prompts (ours) 23.19 (+0.55) 0.707 (+0.009) 46.13 (–7.47) 1096.06 (–233.19)
Hyper-GAE Yang et al. (2023) 22.12 0.682 72.62 1520.49
+ Prompts (ours) 22.46 (+0.34) 0.694 (+0.012) 59.43 (–13.19) 1261.53 (–258.96)
M2DN Meng et al. (2024) 23.47 0.715 42.52 845.29
+ Prompts (ours) 23.84 (+0.37) 0.726 (+0.011) 33.79 (–8.73) 684.62 (–160.67)
C.2 Ablation Study